A metallic metre stick moves with a velocity of 2 m/s in a direction perpendicular to its length and perpendicular to a uniform magnetic field of magnitude 0.2 T. Find the emf induced between the ends of the stick.

Grade:10

1 Answers

Deepak Patra
9 years ago
Sol. ℓ = 1 m, B = 0.2 T, v = 2 m/s, e = Bℓv = 0.2 × 1 × 2 = 0.4 V